40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 17 16:05:07 2011, kiwamu, Update, ABSL, LightWave out of MOPA box DSC_3139_small.png
    Reply  Mon Jun 20 11:38:49 2011, kiwamu, Update, ABSL, I-P curve of LightWave M126-1064-700 IP_curve.png
       Reply  Mon Jun 20 13:48:25 2011, Koji, Update, ABSL, I-P curve of LightWave M126-1064-700 
       Reply  Mon Jun 20 16:44:02 2011, steve, Update, ABSL, I-P curve of LightWave M126-1064-700 
          Reply  Thu Jun 23 22:59:45 2011, kiwamu, Update, ABSL, I-P curve of LWE IP_curve_small.png
             Reply  Fri Jun 24 07:49:23 2011, steve, Update, ABSL, I-P curve of LWE with serial numbers 
Message ID: 4872     Entry time: Thu Jun 23 22:59:45 2011     In reply to: 4842     Reply to this: 4877
Author: kiwamu 
Type: Update 
Category: ABSL 
Subject: I-P curve of LWE 

 The I-P curve was measured again, but this time in a lower current range of 1.0-1.9 [A].

The plot below is the latest I-P curve.

IP_curve_small.png

(Decision)

Based on the measurement and some thoughts, I decided to run this laser at about 1.8 [A] which gives us a middle power of ~ 360 [mW].

In the 40m history, the laser had been driven at 2.4 [A] in years of approximately 2006-2009, so it's possible to run it at such a high power,

but on the other hand Steve suggested to run it with a smaller power such that the laser power doesn't degrade so fast.


(notes)

  The laser controller handed from PK (#4855) was used in this measurement.

The nominal current was tuned to be 1.8 [A] by tuning a potentiometer on the laser head (see page.18 on the manual of LWE).

There was a huge bump around 1.4 [A] and sudden power drop at 1.48 [A] although I don't know the reason.

Quote from #4842

The old days the NPRO ( inside the MOPA ) was running ~1.7A  500 mW

ELOG V3.1.3-